### v2

@@ 3.1 DISCLOSURE CONTROL IMPLEMENTED
For tabular data (e.g., cross tabulation and pivot tables), ACRO prohibits the reporting of the maximum or minimum value in any cell that represents a sub-group of one or more contributors. Moreover, ACRO suppresses, and reports the reason why, the value of an aggregation statistic (mean, median, variance, etc.) for any cell is deemed to be sensitive. The current version of ACRO supports the three most common tests for sensitivity: ensuring the number of contributors is above a frequency threshold, and testing for dominance via $p\%$ and $NK$ rules. In addition, any outputs found to include negative or missing values are automatically flagged for human review since the results of SDC are not well defined in these circumstances.
For tables, ACRO builds a series of suppression masks, which indicate which cells are to be suppressed for each check. A summary outcome table indicating which suppression rule was applied to each cell is presented to the researcher, along with the result of the query.
For regressions such as linear, probit, and logit, the tests verify that the number of residual degrees of freedom exceeds a threshold. There is currently limited support for graphical plots, however ACRO supports the SDC of histograms using the same rules as for frequencies since these fall under the same statbarn category; if a histogram is found to be disclosive and automatic suppression is enabled, the plot is hidden from the researcher. Additionally, survival plots such as Kaplan-Meier are protected using a rounding mitigation strategy.
As noted above, all of these tests and checks are configurable according to the TRE’s risk appetite. The data custodian, e.g., TRE staff member, specifies the parameter values used for the output checks in a YAML configuration file, which is loaded upon ACRO initialisation. The default ACRO parameters are shown in Table LABEL:table:params. Future releases will offer the option to override these on a dataset, or even attribute level.
@@ 3.2 PYTHON PACKAGE
Python is a popular multi-platform language widely used for data analysis and machine learning. PyPI provides a simple and easy to use package management system for distributing open source Python libraries. Pandas [23] and the statsmodels [24] Python libraries are mature, popular, and well-supported packages for data analysis, statistical testing, and statistical data exploration. Pandas is used by more than 55% of all Python users [25]. The use of Python as the primary implementation therefore enables the leveraging of existing expertise and community support with these packages so that the ACRO front-end can be as similar to the API researchers already know and trust, and further facilitates the rapid development of disclosure checking functionality on the back-end. As the PyPI distribution system is simple and allows the use of semantic versioning, it supports a rapid and iterative develop-and-deploy strategy to provide continuing functionality and improvements. The Python ACRO package also depends on Matplotlib [26] to support data visualisations, NumPy [27] as a core numerical matrices package, and PyYAML [28] for YAML configuration file support.
The newest version of ACRO may be installed from PyPI:
\$ pip install acro
@@
Then, to get started, instantiate an ‘ACRO’ object as seen in Figure LABEL:fig:pyinit, which is used to store and maintain outputs. The ‘ACRO’ constructor optionally accepts a string parameter config that can be used to specify an alternative YAML rule configuration file, and a Boolean parameter suppress that specifies whether automatic suppression should be applied within the session.
@@
Standard data analysis operators can then be used by pre-fixing the function with ‘acro’. Figure LABEL:fig:crosstab shows how the ‘pandas.crosstab’ function can be used with the same arguments and in the same way, operating on a ‘pandas.DataFrame’, df. However, in this instance, the R/G column and a single cell in column N is automatically suppressed since it violates disclosure checks. The output also includes an explanation of which disclosure checks failed for each cell, a summary of the checks, and the name assigned to the output so that it can be referenced subsequently; for example to add comments or request an exception from a human checker.
@@
Similarly, pivot tables can be created as in Figure LABEL:fig:pivot. Note that the only API difference is calling ‘acro.pivot_table’ instead of ‘pandas.pivot_table’. In this example, the entire column R/G has been withheld from the researcher as a result of SDC. Since the tables in these examples failed SDC, exception requests must be provided, such as as in Figure LABEL:fig:exception.
@@
In addition to common Pandas tabular operations, ACRO provides disclosure control functions for common statsmodels regression operations. Figure LABEL:fig:ols shows an example where an ordinary least squares is performed on $x$ and $y$ where the residual degrees of freedom are checked. Note that the only API difference is calling ‘acro.ols’ instead of ‘statsmodels.api.OLS’.
@@
Limited support for graphical plots is available. For example, Figure LABEL:fig:hist shows the result of attempting to produce a disclosive histogram. The plot is therefore hidden from the researcher. Note that the only API difference is calling ‘acro.hist’ instead of ‘pandas.hist’.
@@
Moreover, as previously shown the researcher can request an exception for a particular output and include their supporting evidence; unsupported outputs can be added with the add_custom_output function and by specifying the path to any file. These unsupported outputs are then included in the final requested release and highlighted for the attention of a human checker. When the researcher is finished with their analysis, they simply call the finalise function:
$>$ acro.finalise()
@@
The automatic suppression of outputs can be enabled or disabled via the Boolean suppress constructor parameter. Regardless of whether suppression is applied, the results of the checks are included in the final JSON report. This report file can then be ingested by the GUI described in Section 4 to display the disclosure results and suppression can be applied on a case by case basis.
@@
Some Jupyter notebooks [39] demonstrating example code usage and output are available within the ACRO project repository. The currently implemented methods (v0.4.6) are listed as follows; for more details see the ACRO API documentation built with the tool Sphinx [40], which is available: https://ai-sdc.github.io/ACRO/.
@@ 3.3 R PACKAGE
The R ACRO package [4] is an example of cross-language support and is available through CRAN. It provides a set of R wrapper functions that execute Python back-end checking within a virtual environment via the reticulate package [41], which provides automatic conversions for types such as the R ‘data.frame’ to Pandas ‘DataFrame’. The source code is available under MIT license at: https://github.com/AI-SDC/ACRO-R. GitHub CI runners automatically validate the package for CRAN by running R CMD check [42] using different operating systems and Python versions to ensure cross-platform operability. CI runners also generate and publish API documentation using pkgdown [43] and generate test code coverage reports.
For regressions, the common R lm and glm functions were shadowed with equivalent versions implemented as acro_lm and acro_glm, respectively, exploiting the fact that statsmodels already supports R-style specification of formulae. For tabular data, the standard R table command has been shadowed with acro_table to perform cross tabulation analysis. While the dplyr package [44] is commonly used, no simple pivot table functions are provided; instead various combinations of groupby and summarize etc. are used. Therefore, at this stage of development, the Python cross tabulation and pivot table functions were directly interfaced with acro_crosstab and acro_pivot_table.
To install the R ACRO package from CRAN:
$>$ install.packages("acro")
Then to instantiate an ‘ACRO’ object:
@@
Subsequently, the Python package functions may be used, however for simplicity the package provides R helper functions that use underscore naming. For example, the crosstab function can be used identically to the Python version by calling acro_crosstab. Similarly, the pivot_table function becomes acro_pivot_table, etc. See the following list of R helper functions available; see the ACRO R reference manual on CRAN [^5] https://cran.r-project.org/web/packages/acro/acro.pdf for further details.
3.3.1 R TABLE METHODS
3.3.2 R REGRESSION METHODS
3.3.3 R SESSION MANAGEMENT METHODS
@@ 4 GRAPHICAL USER INTERFACE
OpenSAFELY have developed a platform-independent stand-alone tool for output checkers to view the outputs produced by ACRO, understand the risks associated with each output, make audited decisions, and produce zipped packages of files for release.
@@
The application is packaged using standard tooling, e.g., as an msi for Windows, deb for Debian/Ubuntu, and dmg for macOS operating systems; all dependencies are bundled with the installers, including a chrome-based web browser [48], Python, and JavaScript runtimes. The Cypress [49] Javascript framework is used to implement a number of automated user interface tests for CI.
@@
A screenshot of the GUI can be seen in Figure LABEL:fig:gui. The left pane displays the list of outputs to review. The top right pane provides an option to view the ACRO configuration specifying the risk profile. The centre right pane shows details of the currently selected output, including comments from the researcher, any exception requests, and the ACRO recommendation.
Moreover, the GUI provides buttons for the TRE checker to approve or reject the output and supply feedback, e.g., if the checker wishes to override the ACRO recommendation. The bottom right pane displays the currently selected output with any specific components that violate SDC checks highlighted in red. When checking has been completed, the top right ‘Release and download’ button can be used to record the overall comments and create a (zip) package for release, which automatically excludes any rejected outputs.
@@ 5 FUTURE PLANS
Future plans include ongoing development of both the main Python ACRO library and additional translation languages such as R and Stata to enhance functionality, support more analyses and respond to inevitable issues as they arise from an increased volume of testing. Future plans also include the production of an easy to install package for Stata; and the development of a toolkit that can be embedded within projects such as OpenSAFELY and medical TREs.
Future plans include implementing rounding mitigation strategies for frequency tables, improved support for graphical plots, and the ability to configure rules at a dataset or even attribute level.
Currently, the results of SDC checks for tables are presented by displaying the whole table, where the reasons for failure are presented within each cell; while this works well for small to medium size tables, it can be overwhelming for larger tables. Future versions may therefore present the row/column numbers of the offending cells with an explanation of the SDC check instead of the entire table.
Additional features and improved user experience will be facilitated by the further involvement of end-users and output checkers. Future versions may also extend coverage of more of the R Tidyverse [50] family such as dplyer, etc.; and rewriting parts of the translation to support different Stata versions.
As previously mentioned, the current version of ACRO does not implement secondary disclosure checks. Work is underway investigating how to at least flag potential differencing risks across the set of outputs from a research session, and in future, create a library of outputs which might allow secondary disclosure to be assessed, even if only partially.
Moreover, future versions of the GUI viewer may include outputs from checking trained machine learning models, such as those from the sacroml [51] Python package.
As the TRE infrastructure landscape evolves, and new lightweight dashboard frameworks such as Streamlit [52] increase adoption, the question of choice of tools for implementing the viewer design inevitably remain open.
The GUI viewer is currently intended for checkers to review the results produced by researchers, but not for researchers themselves to view, modify, and supply information for checkers; therefore, future plans include the development of such a GUI for researchers.

### v3

@@ 3.2 DISCLOSURE CONTROL IMPLEMENTED
For tabular data (e.g., cross tabulation and pivot tables), the ACRO tool prohibits the reporting of the maximum or minimum value in any cell that represents a sub-group of one or more contributors. Moreover, ACRO reports the reason why the value of an aggregation statistic (mean, median, variance, etc.) for any cell is deemed to be sensitive.
The current version of ACRO supports the three most common tests for sensitivity: ensuring the number of contributors is above a frequency threshold, and testing for dominance via $p\%$ and $NK$ rules. Sensitive cells may optionally be suppressed. In addition, any outputs found to include negative or missing values are automatically flagged for human review since the results of SDC are not well defined in these circumstances.
For tables, ACRO builds a series of masks that indicate which cells fail the sensitivity tests for each check. For reach check there are two stages; first a version of the requested table is built using a custom aggregation function (or ‘count’ for cell thresholds). The resulting tables are then binarised via comparison with a parameter that represents the data-owners’ appetite for that risk. A summary outcome table indicating which rule was applied to each cell is presented to the researcher, along with the result of the query. If suppression is enabled, the offending cells are censored.
For regressions such as linear, probit, and logit, the tests verify that the number of residual degrees of freedom exceeds a threshold. There is currently limited support for graphical plots, however ACRO supports the SDC of histograms using the same rules as for frequencies since these fall under the same statbarn category; if a histogram is found to be disclosive and suppression is enabled, the plot is hidden from the researcher. Additionally, survival plots such as Kaplan-Meier may be protected using a rounding mitigation strategy.
As noted above, all of these tests and checks are configurable according to the TRE’s risk appetite. The data custodian, e.g., TRE staff member, specifies the parameter values used for the output checks in a YAML configuration file, which is loaded upon ACRO initialisation. The default ACRO parameters are shown in Table LABEL:table:params. Future releases will offer the option to configure parameters on a dataset or attribute level, in addition to the current session-based method.
@@ 3.3 PYTHON PACKAGE
Python is a popular multi-platform language widely used for data analysis and machine learning. PyPI provides a simple and easy to use package management system for distributing open source Python libraries. Pandas [25] and the statsmodels [26] Python libraries are mature, popular, and well-supported packages for data analysis, statistical testing, and statistical data exploration. Pandas is used by more than 55% of all Python users [27].
The use of Python as the primary implementation therefore enables the leveraging of existing expertise and community support with these packages so that the ACRO front-end can be as similar to the API researchers already know and trust, and further facilitates the rapid development of disclosure checking functionality on the back-end.
As the PyPI distribution system is simple and allows the use of semantic versioning, it supports a rapid and iterative develop-and-deploy strategy to provide continuing functionality and improvements. The Python ACRO package also depends on Matplotlib [28] to support data visualisations, NumPy [29] as a core numerical matrices package, and PyYAML [30] for YAML configuration file support.
@@
After installing ACRO from PyPI, a session can be started by instantiating an ‘ACRO’ object as seen in Figure LABEL:fig:pyinit, which is used to store and maintain outputs. Each session would typically be contained within a single period of (physical or virtual) access to a TRE. The ‘ACRO’ constructor optionally accepts a string parameter config that can be used to specify an alternative YAML rule configuration file, and a Boolean parameter suppress that specifies whether suppression should be applied within the session.
@@
Standard data analysis operators can then be used by pre-fixing the function with ‘acro’. Figure LABEL:fig:crosstab shows how the ‘pandas.crosstab’ function can be used with the same arguments and in the same way, operating on a ‘pandas.DataFrame’, df. However, in this instance, the R/G column and a single cell in column N is automatically suppressed since it violates disclosure checks and mitigation has been enabled. The output also includes an explanation of which disclosure checks failed for each cell, a summary of the checks, and the name assigned to the output so that it can be referenced subsequently; for example to add comments or request an exception from a human checker.
@@
Similarly, pivot tables can be created as in Figure LABEL:fig:pivot. Note that the only API difference is calling ‘acro.pivot_table’ instead of ‘pandas.pivot_table’. In this example, the entire column R/G has been withheld as a result of enabling suppression. Since the tables in these examples failed SDC, exception requests must be provided, as in Figure LABEL:fig:exception.
A benefit of building ACRO on the vast user-base of Pandas is the plethora of help and guidance available. For example, should a researcher wish to create a table with survey-weighted data, a simple web query provides examples of how this can be done by defining a custom aggregation function, and passing that to the tabulation command. Note, however, that in this case the checks would be performed using the unweighted data.
@@
In addition to common Pandas tabular operations, ACRO provides disclosure control tests for common statsmodels regression operations. Figure LABEL:fig:ols shows an example where an ordinary least squares is performed on $x$ and $y$ where the residual degrees of freedom are checked. Note that the only API difference is calling ‘acro.ols’ instead of ‘statsmodels.api.OLS’.
@@
Limited support for graphical plots is available. For example, Figure LABEL:fig:hist shows the result of attempting to produce a disclosive histogram. In this case, suppression has been enabled and so the plot is therefore hidden from the researcher. Note that the only API difference is calling ‘acro.hist’ instead of ‘pandas.hist’.
@@
Moreover, as previously shown the researcher can request an exception for a particular output and include their supporting evidence; unsupported outputs can be added with the add_custom_output function and by specifying the path to any file. These unsupported outputs are then included in the final requested release and highlighted for the attention of a human checker. When the researcher is finished with their analysis, they simply call the finalise function.
@@
The automatic suppression of outputs can be enabled via the Boolean suppress constructor parameter; by default it is disabled. Regardless of whether suppression is applied, the results of the checks are included in the final JSON report. This report file can then be ingested by the GUI described in Section 4 to display the disclosure results and mitigation techniques can be applied on a case by case basis.
@@
Some Jupyter notebooks [41] demonstrating example code usage and output are available within the ACRO project repository. The currently implemented methods (v0.4.7) are listed as follows; for more details see the ACRO API documentation built with the tool Sphinx [42], which is available: https://ai-sdc.github.io/ACRO/.
@@ 3.4 R PACKAGE
The R ACRO package [6] is an example of cross-language support and is available through CRAN. It provides a set of R wrapper functions that execute Python back-end checking within a virtual environment via the reticulate package [43], which provides automatic conversions for types such as the R ‘data.frame’ to Pandas ‘DataFrame’. The source code is available under MIT license at: https://github.com/AI-SDC/ACRO-R.
GitHub CI runners automatically validate the package for CRAN by running R CMD check [44] using different operating systems and Python versions to ensure cross-platform operability. CI runners also generate and publish API documentation using pkgdown [45] and generate test code coverage reports.
For regressions, the common R lm and glm functions were shadowed with equivalent versions implemented as acro_lm and acro_glm, respectively, exploiting the fact that statsmodels already supports R-style specification of formulae. For tabular data, the standard R table command has been shadowed with acro_table to perform cross tabulation analysis.
While the dplyr package [46] is commonly used, no simple pivot table functions are provided; instead various combinations of groupby and summarize etc. are used. Therefore, at this stage of development, the Python cross tabulation and pivot table functions were directly interfaced with acro_crosstab and acro_pivot_table.
To start a session, instantiate an ‘ACRO’ object:
@@
Subsequently, the Python package functions may be used, however for simplicity the package provides R helper functions that use underscore naming. For example, the crosstab function can be used identically to the Python version by calling acro_crosstab. Similarly, the pivot_table function becomes acro_pivot_table, etc. See the ACRO R reference manual on CRAN for further details. [^5] https://cran.r-project.org/web/packages/acro/acro.pdf
@@ 4 GRAPHICAL USER INTERFACE
OpenSAFELY developed a platform-independent stand-alone tool for checkers to view the outputs produced by ACRO, understand the risks associated with each output, make audited decisions, and produce zipped packages of files for release.
@@
The application is packaged using standard tooling, e.g., as an msi for Windows, deb for Debian/Ubuntu, and dmg for macOS operating systems; all dependencies are bundled with the installers, including a chrome-based web browser, Python, and JavaScript runtimes. The Cypress [50] Javascript framework is used to implement a number of automated user interface tests for CI.
@@
A screenshot of the GUI can be seen in Figure LABEL:fig:gui. The left pane displays the list of outputs to review. The top right pane provides an option to view the ACRO risk profile. The centre right pane shows details of the currently selected output, including comments from the researcher, any exception requests, and the ACRO recommendation.
Moreover, the GUI provides buttons for checkers to approve or reject the output and supply feedback, e.g., if the checker wishes to override the ACRO recommendation. The bottom right pane displays the currently selected output with any specific components that violate SDC checks highlighted in red. Hovering over the tables displays pop-up information detailing which test failed for each cell. When checking has been completed, the top right ‘Release and download’ button can be used to record the overall comments and create a (zip) package for release, which automatically excludes any rejected outputs.
@@ 5 FUTURE PLANS
Future plans include the ongoing development of both the main Python package and additional translation languages to support more analyses and enhance functionality. Improvements will also result from extending the coverage of the R Tidyverse [51] family such as dplyer, and the production of an easy to install package for Stata.
The user experience will be enhanced by facilitating the selective application of disclosure mitigations on a query by query basis to provide even more control over the process; implementing additional mitigation options such as rounding and differential privacy for tables and statistical models; improved support for graphical plots; and the ability to configure rules at a dataset or even attribute level.
Currently, the results of SDC checks for tables are presented by displaying the whole table, where the reasons for failure are presented within each cell; while this works well for small to medium size tables, it can be overwhelming for larger tables. Presenting the row/column numbers of the offending cells with an explanation of the SDC check instead of the entire table will aid comprehension.
Additional parameterised disclosure checks will increase the effectiveness of the toolkit, such as testing for minimum sample sizes in the use of regressions applied to categorical data. As previously mentioned, the current version of the SACRO toolkit does not implement secondary disclosure checks. However, work is underway investigating how to at least flag potential differencing risks across the set of outputs from a research session and create a library of outputs which might allow secondary disclosure to be assessed, even if only partially, but always with the focus on assisting rather than replacing human checkers.
Moreover, future versions of the GUI viewer may include outputs from checking trained machine learning models, such as those from the sacroml [52] Python package. The GUI viewer has been designed for checkers to review the results produced by researchers; the development of a GUI for researchers will enable them to view, modify, and supply information to checkers. Furthermore, as the TRE infrastructure landscape evolves, and new lightweight dashboard frameworks such as Streamlit [53] increase adoption, the question of choice of tools for implementing the viewer design inevitably remain open.
Future work also includes the development of a toolkit that can be embedded within projects such as OpenSAFELY and medical TREs. Additional features and improved user experience will be facilitated by the further involvement of end-users and output checkers.